OpenStudy (anonymous):

how do i right {x: 0

OpenStudy (anonymous):

The given set describes all the values of \(x\) that give you a positive value of \(\cos x\). In terms of quadrants, \(\cos x>0\) in the first and fourth quadrants. In interval notation, \(\cos x>0\) in the interval \(\left(-\dfrac{\pi}{2},\dfrac{\pi}{2}\right)\).